OpenStudy (anonymous):

Three numbers form a geometric sequence. Their sum is 21 and their product is 64. Find the numbers(please show working)

OpenStudy (anonymous):

a,b,c\[S_n=\frac{ a(1-r^n) }{ 1-r }\] \[\frac{ T_2 }{ T_1 }=\frac{ T_3 }{ T_2 }\]

OpenStudy (anonymous):

So from here how do I continue this is what I've tried so far and it's not getting me anywhere: (T2/T1)=(b/a)r and (T3/T2)=(c/b)r

OpenStudy (anonymous):

b/a=c/b b^2=ac

OpenStudy (anonymous):

so we have\[a+b+c=21\] \[abc=64\] \[b^2=ac\] 3 equations 3 unknowns ,solve

hartnn (hartnn):

u can easily find b from abc=64 and b^2=ac

hartnn (hartnn):

try it, if u can't i'll help u

OpenStudy (anonymous):

k, an trying this out now already got b=4

hartnn (hartnn):

thats correct, now u know how u can get a,c ?

OpenStudy (anonymous):

just got it, a=1,b=4,c=16 thanks alot!